# Simplicity Prevails: The Emergence of Generalizable AIGI Detection in Visual Foundation Models

### v1

@@ 3.2 Performance on Standard Benchmarks
Table 4: Generalization on AIGI-Now. Evaluation on 9 open-source and close-source generative models. Accuracy is averaged over real and fake classes. Best results in bold.
@@
3.3 The Collapse of SOTA in the Wild
While standard benchmarks provide a controlled environment, real-world deployment involves diverse, unconstrained data distributions. To evaluate this, we test on four challenging in-the-wild datasets: Chameleon, WildRF, Social-RF, and CommunityAI. The results, summarized in Table 3, reveal a stark contrast. Results. Most specialized detectors suffer a catastrophic performance collapse. Methods like NPR, LaDeDa, and even recent techniques like OMAT degrade to near-random performance, primarily due to a failure to recognize diverse fake samples. The only exception among specialized methods is DDA, which maintains a respectable average accuracy of 85.0%, due to its alignment with the robust VAE decoder patterns shared across latent diffusion models by training on DINOv2 with VAE reconstruct data.
3.4 Generalization to State-of-the-Art Generators
@@
A critical question remains: do these models truly learn generalized forensic concepts, or do they merely memorize patterns seen during pre-training? To investigate this, we conduct a rigorous evaluation on AIGIHolmes and AIGI-Now. Crucially, these datasets contain:
@@
  • Novel Architectures (AIGIHolmes): Including recent Auto-Regressive (AR) models (e.g., LlamaGen, VAR) and Diffusion Transformers (e.g., FLUX, PixArt-XL), whose generation mechanisms differ fundamentally from the UNet-based diffusion models (e.g., SD-v1.4) dominant in older web corpora and common training data.
  • Proprietary Engines (AIGI-Now): Including closed-source commercial APIs (e.g., GPT-4o, Nano Banana, FLUX-Pro), ensuring that their specific training data and artifacts were strictly inaccessible during the pre-training of our VFM backbones.
Results. As summarized in Tables 3.2 and 4, modern VFMs demonstrate exceptional transferability across these unseen domains. On AIGIHolmes, PE-CLIP-Linear and DINOv3-Linear achieve remarkable average accuracies of 97.8% and 97.2%. Notably, they maintain high performance even on Auto-Regressive models (e.g., VAR: 93.5%/92.2%), confirming that their learned features capture fundamental anomalies of synthesis rather than specific architecture-dependent artifacts. Similarly, on AIGI-Now, MetaCLIP2-Linear leads with 90.7%, successfully detecting closed-source generators like GPT-4o. Conversely, specialized detectors struggle significantly, with methods like CNNSpot collapsing to near-random guessing ($\sim$57%) on these novel distributions. This confirms that modern VFMs have emerged with a generalized notion of artificiality that extends well beyond the specific generators seen during pre-training.
4 Analysis: The Mechanisms of Emergence
Our empirical results establish a counter-intuitive finding: simple linear probes on frozen Vision Foundation Models (VFMs) consistently outperform specialized detectors that employ complex, domain-specific architectures. This raises a fundamental question regarding the nature of this capability. Is this an architectural superiority inherent to modern Transformers, or is it an artifact of the training data? In this section, we deconstruct the source of this generalizability and propose that the observed performance is not forensic “intelligence” in the traditional sense, but rather a direct manifestation of massive-scale data exposure.
Figure 1: The Surge of Generative Data in Web Corpora. We track the number of indexed pages from civitai.com within Common Crawl snapshots from 2022 to 2025.
4.1 The Source: An Explosion of Generative Content
To investigate synthetic data exposure in pre-training corpora, we analyzed the Common Crawl (CC) index, the primary data reservoir for models like MetaCLIP and DINOv3. Given the difficulty of quantifying in-the-wild AIGI, we use the indexing volume of civitai.com, the largest open-source generative model platform, as a proxy for the prevalence of high-quality synthetic content.
Figure 1 reveals a distinct phase shift. The pre-2022 “Silent Era” contained negligible generative content, explaining why legacy models like OpenAI CLIP lack forensic capability. In contrast, the period from 2023 to 2025 shows an exponential surge. Despite a transient dip in mid-2025 due to anti-crawling updates, the trajectory remains aggressively upward, peaking at over 40,000 indexed records per snapshot in late 2025. This rise directly correlates with the proliferation of tools like SDXL and FLUX.
This implies that modern VFMs have been implicitly supervised on millions of AIGI samples alongside their textual descriptions. This “contamination” effectively transforms unsupervised pre-training into a weakly-supervised detection task. We next trace how this exposure manifests in Vision-Language versus Self-Supervised Models.
4.2 Mechanism I: Semantic Conceptualization in VLMs
@@
  • Source-Related Concepts: Specific names of generative models or platforms (e.g., ’GenImage’, ’ADM’, ’BigGAN’, ’glide’, ‘Midjourney’).
@@
Table 5: Comparison of Text–Image Similarities on In-the-Wild Dataset
Table 5 presents the results of our semantic probing, revealing a striking dichotomy between legacy and modern VLMs. Legacy models like CLIP (2021) and SigLIP (2023) exhibit ”forensic blindness,” mapping synthetic images to content terms (e.g., portrait). Notably, even the recently released SigLIP 2 (2025) fails to detect forgery concepts (Top-1: genuine/urban), likely because it relies on the older WebLI dataset (Chen et al., 2022) curated in 2022, prior to the generative explosion. In sharp contrast, modern VLMs trained on recent web crawls (MetaCLIP 2, PE) consistently align fake images with “AI generated”. Crucially, on Midjourney-CC, these models specifically retrieve “midjourney_images”, providing definitive evidence that their capability stems from exposure to recent, platform-specific metadata which older datasets lack.
4.3 Mechanism II: Data-Driven Feature Discrimination in SSL
@@
Table 6: Counterfactual Analysis. Comparison of DINOv3 trained on Web Data vs. Satellite Data. The satellite model suffers a catastrophic collapse on in-the-wild fakes.
@@
5 Robustness and Limitations
5.1 Protocol I: Resilience to Common Perturbations
We conduct our evaluation on the Chameleon dataset, which represents the most challenging in-the-wild distribution. We benchmark against DDA, the only specialized detector that maintains non-trivial performance on this dataset. We apply varying levels of JPEG compression (Quality $\in\{95,\dots,65\}$) and Gaussian Blur ($\sigma\in\{0.5,\dots,2.0\}$) to the test set.
Table 7: Robustness to Common Perturbations on Chameleon Dataset. We evaluate the impact of JPEG compression and Gaussian Blur.
Detector Base JPEG Compression (Quality $\downarrow$) Gaussian Blur ($\sigma\uparrow$)
(No Aug) 95 85 75 65 0.5 1.0 1.5 2.0
MetaCLIP-Linear 0.644 0.614 0.667 0.614 0.749 0.614 0.744 0.775 0.722
SigLIP-Linear 0.606 0.583 0.600 0.593 0.605 0.500 0.477 0.482 0.471
DINOv2-Linear 0.608 0.606 0.605 0.606 0.596 0.590 0.564 0.564 0.542
DDA 0.824 0.836 0.823 0.790 0.790 0.844 0.843 0.797 0.758
SigLIP2-Linear 0.858 0.857 0.856 0.860 0.828 0.836 0.776 0.689 0.671
MetaCLIP2-Linear 0.930 0.933 0.940 0.937 0.898 0.922 0.931 0.939 0.932
PE-CLIP-Linear 0.959 0.967 0.964 0.959 0.921 0.938 0.925 0.831 0.778
DINOv3-Linear 0.914 0.918 0.914 0.916 0.891 0.916 0.909 0.897 0.891
As detailed in Table 7, the results reveal a critical trade-off between peak performance and stability among modern VFMs. JPEG compression, which typically disrupts the specific artifact patterns used by traditional detectors, has a minimal impact on our baselines. This confirms that these models do not rely on fragile high-frequency noise statistics but rather on robust feature representations.
A significant divergence appears under Gaussian Blur. PE-CLIP, despite establishing the highest baseline accuracy (95.9%), exhibits a sharp performance degradation as blur intensity increases, dropping to 77.8% at $\sigma=2.0$. This indicates that PE-CLIP’s superior detection capability is partially driven by fine-grained, high-frequency textual or visual artifacts that are easily erased by smoothing. In sharp contrast, MetaCLIP2 and DINOv3 demonstrate remarkable resilience. MetaCLIP2 actually maintains or slightly improves its accuracy under blur, while DINOv3 experiences a negligible drop. This indicates that the models capture fundamental structural anomalies inherent to the synthesis process, which are robust to smoothing operations, unlike local pixel-level artifacts. This property makes them ideal candidates for real-world applications involving low-resolution previews or video frame analysis. We also provide additional experiments verifying their efficacy on video detection benchmarks in Appendix B.
5.2 Protocol II: Real-World Transmission and Recapture
@@
Table 8: Robustness Evaluation on RRDataset. We report the accuracy on Real and AI classes separately.
@@
5.3 Protocol III: Robustness to Reconstruction and Editing
@@
Table 9: Limitations under Reconstruction and Editing. Detection accuracy on DDA-COCO (VAE-based reconstruction) and BR-Gen (Diffusion-based local editing).
@@
6 Conclusion
In this work, we revisited AIGI detection through the lens of the “Bitter Lesson,” demonstrating that simplicity prevails over complex architectural designs. We show that simple linear probes on modern Vision Foundation Models (VFMs) achieve state-of-the-art performance, decisively outperforming specialized detectors in challenging in-the-wild scenarios.Our analysis traces this capability to pre-training data exposure rather than architectural bias. Modern VFMs have effectively internalized the ”generative manifold” from web-scale data, either through semantic alignment (VLMs) or implicit feature discrimination (SSL).However, we also identify the boundaries of this paradigm: while VFMs solve the generalization problem for fully synthetic content, they remain blind to fine-grained reconstruction and localized editing.We conclude that the future of AI forensics lies not in hand-crafting artifact extractors, but in harnessing the evolving world knowledge of foundation models. As the line between real and synthetic blurs, leveraging the very models driving this revolution offers the most robust path forward.
Appendix A Survey of Backbones in Recent AIGI Detection Methods
To contextualize our work within the current literature, we provide a comprehensive survey of the backbones utilized by state-of-the-art AIGI detection methods published between 2023 and 2025. As shown in Table 10, the vast majority of recent methods rely on legacy architectures, primarily the original CLIP ViT-L/14 (released in early 2021) or standard ResNet-50. Only a few recent works (e.g., DDA) have begun to explore newer self-supervised models like DINOv2. This highlights a significant lag in the forensics community in adopting the latest generation of foundation models, further motivating our investigation into modern VFMs like MetaCLIP 2, SigLIP 2, and DINOv3.
Table 10: Backbone Architectures in Recent AIGI Detection Methods (2023–2025). The table surveys representative methods from top-tier venues. Most existing works are built upon legacy backbones such as OpenAI CLIP (2021) or ResNet-50, lacking exposure to the modern generative web distribution.
Method Year Venue Key Backbone(s) Note
UnivFD (Ojha et al., 2023) 2023 CVPR CLIP-ViT (ViT-L/14) Pioneered CLIP for forensics
DE-FAKE (Sha et al., 2023) 2023 CCS CLIP-ViT (ViT-L/14), ResNet-18 Multi-modal features
FatFormer (Liu et al., 2024) 2024 CVPR CLIP-ViT (ViT-L/14) Adaptive Transformer
NPR (Tan et al., 2024b) 2024 CVPR ResNet-18 (Lightweight) Neighboring Pixel Reconstruction
RINE (Koutlis and Papadopoulos, 2024) 2024 ECCV CLIP ViT-L/14, ViT-B/32 Intermediate representations
CoDE (Baraldi et al., 2024) 2024 ECCV ViT-Tiny Contrastive learning
DRCT (Chen et al., 2024a) 2024 ICML ConvNeXt-Base, CLIP-ViT (ViT-L/14) Reconstruction discrepancy
Patch Shuffle (Zheng et al., 2024) 2024 NeurIPS ResNet-50 Breaking semantic artifacts
C2P-CLIP (Tan et al., 2025) 2025 AAAI CLIP-ViT (ViT-L/14) Prompt tuning
SAFE (Li et al., 2025b) 2025 KDD ResNet-50 (Lightweight) Image transformation
AIDE (Yan et al., 2024a) 2025 ICLR ResNet-50, OpenCLIP (ConvNeXt) Frequency analysis
Effort (Yan et al., 2024b) 2025 ICML CLIP-ViT (ViT-L/14) Orthogonal decomposition
D3 (Yang et al., 2025) 2025 CVPR CLIP-ViT (ViT-L/14) Discrepancy learning
DDA (Chen et al., 2025b) 2025 NeurIPS DINOv2 Dual Data Alignment
OMAT (Zhou et al., 2025a) 2025 NeurIPS CLIP-ViT (ViT-L/14) Latent prior bias
WaRPAD (Choi et al., 2025) 2025 NeurIPS DINOv2 Cropping Robustness
Appendix B Generalization to Video Forgery Detection.
To further evaluate the versatility of the advanced VFM paradigm, we extend our analysis to the temporal domain. We hypothesize that the robust feature representations learned by the latest Vision Foundation Models (VFMs) are not merely restricted to static artifacts but possess inherent cross-task generalization capabilities for video detection.
Methodology and Experimental Setup. For video inference, we employ a straightforward temporal aggregation strategy. Given an input video sequence, we extract a continuous segment of $S$ frames (where $S_{max}=8$). Each frame is processed independently through the frozen VFM backbone to generate frame-level logits. The final prediction for the entire video sequence is derived by calculating the mean of the aggregated logits across all $S$ frames. This approach allows us to leverage the VFM’s ‘firepower’ without the need for specialized temporal architectures or costly fine-tuning on video-specific datasets.
We conducted benchmarks on two prominent datasets: VidproM and GenVideo. Performance is measured using the Recall metric for synthetic subsets (e.g., LaVie, Hotshot, Sora) to assess detection sensitivity.
Table 11: Cross-task Generalization to Video Forgery Detection (VidProm).
Accuracy ↑ (%) Method Pika VC2 T2VZ MS Avg CNNSpot (Wang et al., 2020) 0.5117 0.5018 0.4997 0.5031 0.5041 FreDect (Doloriel and Cheung, 2024) 0.5007 0.5403 0.6988 0.6994 0.6098 Fusing (Ju et al., 2022) 0.5060 0.5007 0.4981 0.5128 0.5044 Gram-Net (Liu et al., 2020) 0.8419 0.6742 0.5248 0.5046 0.6364 GIA (Frank et al., 2020) 0.5373 0.5175 0.4105 0.6022 0.5169 UnivFD (Ojha et al., 2023) 0.4941 0.4865 0.4958 0.5743 0.5127 ReStraV (Internò et al., 2025) 0.9090 0.9950 0.9905 0.9837 0.9706 CLIP 0.9637 0.7764 0.8535 0.1093 0.6757 SigLIP 0.3867 0.9057 0.9987 0.8278 0.7797 SigLIP2 0.7540 0.9896 0.9967 0.9928 0.9333 PE 0.9857 0.9992 0.9926 0.8115 0.9473 MetaCLIP-v2 0.9784 0.9982 0.9971 0.8762 0.9625 DinoV3 0.9985 0.9988 0.9764 0.9768 0.9876
Table 12: Cross-task Generalization to Video Forgery Detection (GenVideo).
Model Testing Subset Avg.
Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild
Studio Valley Scope Scrape
NPR (Tan et al., 2024b) 0.5536 0.7757 0.7188 0.4860 0.0721 0.0429 0.8626 0.6029 0.7143 0.3153 0.5144
STIL (Gu et al., 2021) 0.7500 0.7943 0.9449 0.5786 0.5314 0.6414 0.9712 0.8529 0.6943 0.6242 0.7383
MINTIME-CLIP-B (Coccomini et al., 2024) 0.1607 0.2686 0.1471 0.0557 0.3757 0.0900 0.2188 0.4071 0.0114 0.4309 0.2166
FTCN-CLIP-B (Zheng et al., 2021) 0.9107 0.9300 0.9232 0.1771 0.1750 0.0643 0.9968 0.8479 0.7457 0.5173 0.6287
TALL (Xu et al., 2023b) 0.6071 0.7043 0.9449 0.5914 0.5757 0.6400 0.9633 0.8871 0.6057 0.6210 0.7141
XCLIP-B-FT (Ni et al., 2022) 0.6786 0.9129 0.9623 0.1200 0.2236 0.0914 0.9984 0.8343 0.7557 0.5184 0.6096
DeMamba-XCLIP-FT (Chen et al., 2024b) 0.9286 0.9729 0.9848 0.3829 0.5350 0.4143 0.9984 0.9407 0.7729 0.6415 0.7572
NSG-VD (Zhang et al., 2025) 0.7857 0.9833 0.8000 0.9250 0.9417 0.8750 1.0000 0.9833 0.6833 0.8250 0.8802
Skyra-RL-GenVideo(7B) (Li et al., 2025c) 0.9564 0.9586 0.9623 0.8271 0.9429 0.7200 0.9792 0.9807 0.6629 0.7863 0.8766
CLIP 0.5857 0.9200 0.8942 0.3471 0.6293 0.3100 0.9840 0.9886 0.8557 0.6090 0.7124
SigLIP 0.5000 0.5771 0.3949 0.8429 0.7921 0.7729 0.3626 0.6767 0.2571 0.6982 0.5875
SigLIP2 0.3214 0.9314 0.9036 0.6914 0.7971 0.8600 0.9968 0.9850 0.8129 0.7258 0.8025
PE 0.4286 0.9657 0.9884 0.8114 0.8936 0.7900 1.0000 0.9964 0.9543 0.8216 0.8650
MetaCLIP-v2 0.6871 0.9914 0.9884 0.8300 0.9036 0.8714 1.0000 1.0000 0.9543 0.7148 0.8941
DinoV3 0.6250 0.9957 0.9928 0.9214 0.9571 0.9500 1.0000 0.9971 0.9743 0.7302 0.9144
Observations on Scaling and Quality. The results, summarized in Tab LABEL:tab:VFM-videoBench, reveal a clear correlation between the ‘scale-up’ of VFM parameters/data and their zero-shot detection efficacy. For lower-quality video generation models like LaVie and Hotshot, there is a dramatic performance leap across VFM generations. For instance, while a standard CLIP backbone achieves a mediocre Recall of 62.93% on LaVie, the SigLIP model pushes this to 79.21%. Remarkably, latest-generation VFMs such as DINOv3 and MetaCLIP-2 decisively dominate these benchmarks with Recall rates exceeding 90%.
Interestingly, this performance gain remains evident, albeit less pronounced, when faced with high-fidelity outputs from advanced generators like Sora. Even as the visual quality of forgeries improves, the detection capability of the latest VFMs continues to scale linearly with the model’s pre-training breadth. Perhaps most significantly, despite the absence of video-specific training or official temporal supervision, these general-purpose ‘guns’ consistently outperform bespoke AI video detectors specifically designed for these tasks.
Appendix C Model Architectures and Training Details
In this section, we provide detailed specifications for the Vision Foundation Model (VFM) backbones employed in our study. As summarized in Table 13, we selected a diverse set of state-of-the-art architectures to benchmark the “Simplicity Prevails” hypothesis.
Backbone Selection. Our evaluation suite encompasses two distinct pre-training paradigms:
  • Vision-Language Models (VLMs): Including the MetaCLIP and SigLIP families, alongside the video-enhanced Perception Encoder (PE). These models are pre-trained on billion-scale image-text pairs (e.g., WebLI, Common Crawl), equipping them with the semantic understanding required for the zero-shot concept alignment observed in our experiments.
  • Self-Supervised Models (SSL): Specifically the DINO family. Notably, we utilize the massive DINOv3 (ViT-7B architecture) trained on the LVD-1689M dataset. This uncurated visual exposure allows the model to learn structural forensic features without textual supervision.
Implementation. Consistent with our protocol, all backbones are frozen during training. We attach a simple linear layer to the feature dimension specified in the table. Input images are resized to the native resolution of each model (e.g., $384\times 384$ for SigLIP models) to ensure optimal feature extraction.
Table 13: Backbone Models and Training Specifications. We detail the specific architectures, input resolutions, and pre-training data sources for all VFM baselines evaluated in this work.
Detector Backbone Family Specific Model Feature Dim Input Size Training Data
Vision Transformers (CLIP-based)
MetaCLIP-Linear CLIP MetaCLIP-H/14-2.5B 1280 $\times224$ MetaCLIP 400M
MetaCLIP2-Linear CLIP MetaCLIP-2 Worldwide Giant 1664 $\times224$ Common Crawl (Curated)
SigLIP-Linear CLIP SigLIP-Large/16 1024 $\times384$ WebLI (Chen et al., 2022)
SigLIP2-Linear CLIP SigLIP-2 Giant/16 1536 $\times384$ WebLI (Chen et al., 2022)
PE-CLIP-Linear CLIP PE-Core-L/14 1024 $\times336$ MetaCLIP Images + 22M Videos
Vision Transformers (Self-supervised)
DINOv2-Linear DINO DINOv2-giant 1536 $\times224$ LVD-142M
DINOv3-Linear DINO DINOv3-ViT-7B/16 1664 $\times224$ LVD-1689M
Notes: (1) CLIP-based models are trained on image-text pairs; DINO models use self-distillation. (2) Data sources: WebLI (Chen et al., 2022); LVD datasets (Oquab et al., 2023). (3) All detectors use frozen backbones with linear heads trained on the GenImage dataset.

### v2

@@ 3.3. The Collapse of SOTA in the Wild
Table 4. Generalization on AIGI-Now. Evaluation on 9 open-source and closed-source generative models. Accuracy is averaged over real and fake classes. Best results in bold.
@@
3.4. Generalization to State-of-the-Art Generators
@@
A critical question remains: do these models truly learn generalized forensic concepts, or do they merely memorize pre-training patterns? We investigate this via AIGIHolmes and AIGI-Now, which challenge detectors across two fronts:
@@
  • AIGIHolmes: Features recent Auto-Regressive (AR) models (e.g., LlamaGen, VAR) and Diffusion Transformers (e.g., FLUX), whose mechanisms differ fundamentally from older UNet-based models (e.g., SD-v1.4).
  • AIGI-Now: Contains closed-source APIs (e.g., GPT-4o, FLUX-Pro) unseen during VFM pre-training. And disentangles evaluation into two subsets: Pixel-artifact (pix), which isolates low-level generative traces by strictly aligning image formats; and Semantic (sem), which applies aggressive degradations to obliterate low-level artifacts, forcing detectors to rely solely on high-level semantic anomalies.
Results. As shown in Tables 3 and 4, modern VFMs demonstrate exceptional transferability. On AIGIHolmes, PE-CLIP and DINOv3 achieve average accuracies of 97.8% and 97.2%, maintaining robust detection even on fundamentally distinct AR models (e.g., VAR).
On AIGI-Now, MetaCLIP2 leads with 90.7%. Crucially, VFMs excel across both the pix and heavily degraded sem subsets, proving they capture a dual-level (structural and semantic) notion of artificiality. Conversely, specialized detectors like CNNSpot collapse to near-random guessing ($\sim$50%) on the sem splits. This confirms that modern VFMs learn generalized, robust forensic concepts that extend far beyond specific generators or brittle low-level artifacts.
Table 5. Comparison of Text–Image Similarities on In-the-Wild Dataset
4. Analysis: The Mechanisms of Emergence
The strong performance of simple probes on modern VFMs raises a central question: does this capability primarily come from forensic-specific architectural choices, or from properties already present in large-scale pretrained representations? Motivated by the rapid growth of synthetic content on the web (Figure 1), we investigate whether pre-training data exposure is an important factor behind this phenomenon. Because fully controlled retraining of proprietary billion-parameter models is computationally infeasible, we do not attempt to make a definitive causal claim. Instead, we use a set of complementary indirect analyses to characterize how this capability may emerge. Taken together, these analyses suggest two main mechanisms: semantic conceptualization in Vision-Language Models and implicit distribution discrimination in Self-Supervised Learning models.
4.1. Mechanism I: Semantic Conceptualization in VLMs
@@
  • Source-Related Concepts: Specific names of generative models or platforms (e.g., ‘GenImage’, ‘ADM’, ‘BigGAN’, ‘glide’, ‘Midjourney’).
@@
Table 5 presents the results of our semantic probing, revealing a striking dichotomy between legacy and modern VLMs. Legacy models like CLIP (2021) and SigLIP (2023) exhibit “forensic blindness”, mapping synthetic images to content terms (e.g., portrait). Notably, even the recently released SigLIP 2 (2025) fails to detect forgery concepts (Top-1: genuine/urban), likely because it relies on the older WebLI dataset (Chen et al., 2022) curated in 2022, prior to the generative explosion. In sharp contrast, modern VLMs trained on recent web crawls (MetaCLIP 2, PE) consistently align fake images with “AI generated”. Crucially, on Midjourney-CC, these models specifically retrieve “midjourney_images”, providing definitive evidence that their capability stems from exposure to recent, platform-specific metadata which older datasets lack.
4.2. Mechanism II: Data-Driven Feature Discrimination in SSL
@@
Table 6. Counterfactual Analysis. Comparison of DINOv3 trained on Web Data vs. Satellite Data.
@@
5. Robustness and Limitations
5.1. Protocol I: Resilience to Common Perturbations
To assess real-world reliability, we evaluate detector robustness against JPEG compression (Quality $\in\{95,\dots,65\}$) and Gaussian Blur ($\sigma\in\{0.5,\dots,2.0\}$). We benchmark modern VFM linear probes against legacy models and specialized detectors across both GenImage and the in-the-wild Chameleon datasets.
Figure 2. Robustness to Common Perturbations. Accuracy trajectories under JPEG compression and Gaussian Blur.
As visualized in Figure 2, the trajectories reveal a stark stratification between legacy VFMs (dashed lines) and modern VFMs (solid lines). Older models not only establish lower baselines but also exhibit severe volatility under perturbations, particularly on Chameleon. This decisively proves that robustness is not an inherent benefit of the linear probing architecture. Instead, modern VFMs exhibit superior resilience because their intrinsic forensic capabilities are derived from massive, inadvertent exposure to diverse synthetic content during web-scale pre-training. By internalizing the generative manifold directly from the messy, unconstrained internet, these models learn robust, high-level features rather than brittle, lab-generated artifacts. While PE still relies somewhat on high-frequency traces susceptible to low-pass filtering (dropping to 77.8% at $\sigma=2.0$), DINOv3 and MetaCLIP2 capture fundamental structural anomalies that inherently resist smoothing. This profound resilience to blur and compression allows our image-trained linear probes (e.g., DINOv3) to achieve generalized SOTA performance on video benchmarks like VidProM and GenVideo via simple frame-level aggregation, decisively outperforming bespoke AI video detectors.
5.2. Protocol II: Real-World Transmission and Recapture
@@
Table 7. Robustness Evaluation on RRDataset. We report the accuracy on Real and AI classes separately.
@@
5.3. Protocol III: Robustness to Reconstruction and Editing
@@
Table 8. Limitations under Reconstruction and Editing. Detection accuracy on DDA-COCO (VAE-based reconstruction) and BR-Gen (Diffusion-based local editing).
@@
5.4. Protocol IV: The Bottleneck of Specialized Architectures
A natural question arises from our primary finding: if modern Vision Foundation Models (VFMs) possess such powerful representations, could their performance be further amplified by attaching State-Of-The-Art specialized forensic architectures? To investigate this, we upgraded recent expert models—Effort (Yan et al., 2024b), AIDE (Yan et al., 2024a), and DDA (Chen et al., 2025b)—by swapping their original legacy backbones with our top-performing modern VFMs (MetaCLIP2, PE, and DINOv3).
Table 9. Impact of Upgrading Specialized Architectures. While modern backbones improve the performance of specialized methods, they still severely underperform the simple frozen linear probe. Original baselines are in italics, best results in bold.
Method & Backbone GenImage Chameleon AIGIHolmes
Effort (Original: CLIP-L) 0.807 0.588 0.722
Effort + MetaCLIP2 0.748 0.720 0.799
Effort + PE 0.856 0.761 0.924
AIDE (Original: ResNet/ConvNeXt) 0.768 0.574 0.970
AIDE + MetaCLIP2 0.834 0.709 0.922
AIDE + PE 0.883 0.914 0.947
DDA (Original: DINOv2) 0.890 0.824 0.963
DDA + MetaCLIP2 0.557 0.563 0.596
DDA + PE 0.534 0.559 0.659
DDA + DINOv3 0.572 0.751 0.713
MetaCLIP2-Linear 0.892 0.930 0.942
PE-CLIP-Linear 0.938 0.959 0.978
DINOv3-Linear 0.964 0.914 0.972
The results in Table 9 expose the Bottleneck of Inductive Bias. While upgrading expert models (e.g., AIDE, Effort) with modern VFMs improves their baseline performance, they still strictly underperform our minimalist linear probe (e.g., AIDE with PE achieves 91.4% on Chameleon, falling short of PE-Linear’s 95.9%). Worse still, rigidly specialized architectures like DDA suffer catastrophic degradation ($\sim$50-60%) when forced onto new generic feature spaces. This reveals that complex inductive biases—such as explicit frequency filtering or strict VAE alignment—actually act as information bottlenecks, inadvertently constraining the raw, universal discriminative power naturally emergent in modern representations.
5.5. Protocol V: The Pitfall of Parameter-Efficient Fine-Tuning
Another prevalent paradigm for adapting foundation models is Parameter-Efficient Fine-Tuning (PEFT). If a simple linear layer suffices, would unfreezing the backbone via Low-Rank Adaptation (LoRA) yield even better task-specific performance? To test this, we applied LoRA with rank $r\in\{4,8\}$ to modern VFMs, fine-tuning them on the GenImage (SD v1.4) training set.
Table 10. LoRA Fine-Tuning vs. Frozen Linear Probe. Unfreezing the backbone via LoRA significantly degrades generalization in the wild. Best results in bold.
Backbone Strategy GenImage Chameleon AIGIHolmes
MetaCLIP2 Linear Probe (Frozen) 0.892 0.930 0.942
LoRA (r=4) 0.734 0.817 0.823
LoRA (r=8) 0.780 0.880 0.896
PE Linear Probe (Frozen) 0.938 0.959 0.978
LoRA (r=4) 0.810 0.719 0.879
LoRA (r=8) 0.761 0.635 0.891
DINOv3 Linear Probe (Frozen) 0.964 0.914 0.972
LoRA (r=4) 0.954 0.803 0.977
LoRA (r=8) 0.928 0.718 0.945
Table 10 highlights the Risk of Modifying Internal Knowledge. Contrary to the intuition that fine-tuning improves task adaptation, applying LoRA actively dismantles generalizability. For instance, LoRA fine-tuning on PE (r=8) causes its in-the-wild detection accuracy (Chameleon) to plummet from 95.9% to a mere 63.5%. We attribute this to manifold distortion and catastrophic forgetting. By unfreezing the backbone to optimize for a single generator (SD v1.4), the model rapidly overfits to a narrow, transient distribution of local artifacts, irrevocably overwriting the broad ”world knowledge” of synthetic anomalies it implicitly derived from massive pre-training data.
6. Conclusion
In this work, we revisit AIGI detection from the perspective of pretrained visual representations. We show that frozen features from modern Vision Foundation Models, combined with a lightweight classifier, form a remarkably strong baseline for generalizable AIGI detection. Across standard benchmarks, in-the-wild datasets, and recent unseen generators, this simple setup consistently matches or outperforms recent specialized detectors. Our analyses further suggest that this capability is closely related to exposure to synthetic web content during pre-training, rather than primarily to forensic-specific architectural design. In VLMs, this appears as semantic alignment with forgery-related concepts, while in SSL models it appears as implicit discrimination of generative distributions. Although fully controlled pre-training ablations are beyond the scope of this work, our evidence consistently supports this interpretation.
At the same time, modern VFMs remain weak on pure VAE reconstruction, localized editing, and severe transmission or recapture. We therefore view frozen modern VFM representations not as a complete solution to multimedia forensics, but as a strong foundation for robust global AIGI detection. More broadly, our findings suggest that future progress may depend less on increasingly specialized detector design, and more on effectively leveraging the evolving representations learned by foundation models.